A proton is released from rest in a uniform horizontal electric field. It travels 3.25 m for 5 μs. Find the acceleration of the proton and the magnitude of the electric field.


1
Expert's answer
2021-04-20T11:00:21-0400

To be given in question distance (s)=3.25meter

t=5"\mu sec"

M"=m_{p}"

"q_{p}=q_{e}"

"u=0"

To be asked in question

acceleration (a)=?

Electric force ( F).=?

We know that

Newton's motion 2nd law

"S=ut+\frac{1}{2}at^2"

"3.25=0+\frac{1}{2}a\times(5\times10^{-6})^2"

"a=" 2.6"\times10^{11}meter/sec^2"

Electric force

We know that

"qE =ma"

"E=\frac{m_{p}a}{q_{p}}"

Put value

"E =\frac{1.67\times10^{-27}\times2.6\times10^{11}}{1.6\times10^{-19}}"

"E=2.71375 \times10^{3}N/C"